OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x − y = 6 −4x − 4y = 10

OpenStudy (anonymous):

you must do y=0, so is going to be x-0=6 x=6 replace x=6 in the 2nd question -4*6-4y=10 -10-4y=10 -4y=10+10 y=20/4 y=5